OpenStudy (anonymous):

1/3(9x-6)= 2x+3 what is x=?

OpenStudy (anonymous):

medal

OpenStudy (anonymous):

I'll help

OpenStudy (anonymous):

oh thank you so much

OpenStudy (anonymous):

first distribute the 1/3 into 9 and 6

OpenStudy (anonymous):

??

OpenStudy (anonymous):

3x-2=2x+3

OpenStudy (anonymous):

ok then what?

OpenStudy (anonymous):

now subtract 2x from 3x on both sides

OpenStudy (anonymous):

x-2=3

OpenStudy (anonymous):

add 2 on both sides

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

thanx so much

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i have antother can u help

OpenStudy (anonymous):

@LGumboymath234

OpenStudy (anonymous):

-2n-2/3 = 12 n=?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

add 2/3 on both sides

OpenStudy (anonymous):

-2n=12 2/3

OpenStudy (anonymous):

divide -2 on both sides

OpenStudy (anonymous):

n=-6 1/3